# Redmond, WA %TODO␣FINAL:␣Replace␣with␣your␣institution␣list.www.retrocausal.ai

### v5

@@ 4.1.3 Robustness against Missing Keypoints
To assess the robustness of our method, we perform an experiment where a certain part of the skeleton is removed. Specifically, we randomly remove one finger with a probability $p$ for each frame in the H2O dataset during testing (without any finetuning) and measure the impact of this perturbation on the classification accuracy. The results in Tab. 3 show that LA2DS exhibits a high level of robustness to missing keypoints. For example, when one finger is dropped per frame (i.e., $p=100\%$), the decrease in accuracy for LA2DS is minimal (about $1\%$), whereas for CASA [41], the drop is around $8\%$. This is likely because our method represents each joint by a Gaussian distribution and performs self-attention in the spatial domain to leverage dependencies between joints in the same skeleton, whereas CASA [41] models each joint merely by its coordinates and does not conduct spatial self-attention.
@@ 4.1.4 Robustness against Noisy Keypoints
To examine the generalization of our approach to noisy keypoints, we train and test on skeletons of different qualities. In particular, we use ground truth keypoints provided by the H2O dataset, namely clean keypoints, as well as add random noise to them, yielding noisy keypoints. We measure the impact of this perturbation on the classification accuracy. The results in Tab. 4 show that LA2DS exhibits a high level of robustness against noisy keypoints. For instance, when noisy keypoints are used for both training and testing, the accuracy of CASA [41] is decreased by more than 5%, whereas our method has a drop of less than 2%. Similar to the prior section, our method is more robust to noisy keypoints than CASA [41], thanks to our heatmap representation and self-attention in the spatial domain.
@@ 4.2 Comparison Results
Below we quantitatively compare our approach with previous methods in various fine-grained human activity understanding tasks on three public datasets. In addition, some qualitative comparisons are provided in the supplementary material.
@@ 4.3 Fusion Results
Additional Results and Discussions. Due to space constraints, we provide additional results and discussions in the supplementary material. They include our results with different method parts (2D heatmaps, 2D augmentations, spatial transformer, and fusion), which analyze sources of performance gains, our results with estimated 2D poses (computed by OpenPose [8]), which are generally better than our results with projected 2D poses in the above sections and further advance the state-of-the-art in this topic, and discussions of limitations of our approach, including missing depth and context details.
@@ 5 Conclusion
We present a novel 2D skeleton-based self-supervised temporal video alignment framework for fine-grained human activity understanding. In particular, we propose to take 2D skeleton heatmaps as input and employ a video transformer to perform self-attention both in the spatial and temporal domains. This is in contrast with the prior work of CASA, which directly processes 3D skeleton coordinates and does not conduct self-attention in the spatial domain. In addition, for self-supervised learning, we exploit simple heatmap augmentation techniques based on 2D skeletons. Our approach achieves not only better accuracy but also higher robustness than CASA. Moreover, we show superior performance over previous methods on several fine-grained human activity understanding tasks on Penn Action, IKEA ASM, and H2O. Finally, our multi-modality version establishes the state-of-the-art on all metrics and datasets. To our best knowledge, our work is the first to exploit 2D skeleton heatmap inputs and the first to explore multi-modality fusion for temporal video alignment. Our future work will explore learning by temporally aligning videos with extreme viewpoint gaps and fusing with other modalities such as optical flows.
@@ Appendix 0.A Supplementary Material
In this supplementary material, we first provide the details of our implementation in Sec. 0.A.1. Next, the ablation results of different augmentations are included in Sec. 0.A.2, while our complete multi-modality fusion results on all metrics and datasets are presented in Sec. 0.A.3. We then report the ablation results of different method parts in Sec. 0.A.4. The ablation results of different 2D poses are included in Sec. 0.A.5, while our approach is compared with the concurrent works of VSP [78] and STEPs [62] in Sec. 0.A.6. Lastly, we show some qualitative results in Sec. 0.A.7 and discuss the limitations and societal impacts of our work in Sec. 0.A.8.
@@ 0.A.3 Multi-Modality Fusion
Our complete multi-modality fusion results on all metrics and datasets are presented in Tab. 11. It is clear from the results, our RGB+Pose model yields the best numbers across all metrics and datasets, outperforming all the single-modality methods and the multi-modality baseline extended from CASA [41] and establishing the state-of-the-art across all metrics and datasets. Next, our Pose-only model obtains better results than our RGB-only model on Penn Action and H2O, while having generally worse results on IKEA ASM. It likely implies that pose cues are more important on Penn Action and H2O, while RGB cues are more crucial on IKEA ASM. Finally, both our RGB+Pose model and the CASA-based multi-modality baseline outperform their single-modality counterparts respectively, which validates the benefit of multi-modality inputs for fine-grained human activity understanding tasks.
@@ 0.A.5 Effects of Different 2D Poses
0.A.6 Comparisons with VSP [78] and STEPs [62]
In this section, we compare our approach with the concurrent works of VSP [78] and STEPs [62]. VSP [78] uses video inputs, while STEPs [62] employs videos and optical flows. For our approach, we use OpenPose 2D poses (instead of our projected 2D poses). The results on the IKEA ASM dataset are shown in Tab. 14. It can be seen from Tab. 14 that our method performs on par with VSP [78] (i.e., we obtain worse classification results but better retrieval results) and STEPs [62] (i.e., we achieve better Acc@0.1 but worse Acc@0.5 and Acc@1.0).
Method Input Acc@0.1 Acc@0.5 Acc@1.0 Progress $\tau$ AP@5 AP@10 AP@15
IKEA ASM VSP [78] $\square$ 45.00 45.97 47.52 - - 30.23 - -
STEPs [62] $\approx$ + $\square$ 28.59 36.25 37.02 - - - - -
LA2DS (Ours) $\star$ 26.78 32.79 34.41 - - 33.49 33.04 32.87
LA2DS (Ours) $\star$ + $\square$ 32.13 34.76 36.33 - - 33.51 33.07 32.97
Table 14: Comparisons with VSP [78] and STEPs [62]. Note that $\square$ denotes direct RGB inputs, $\approx$ denotes optical flow inputs, $\blacktriangle$ denotes 3D pose coordinate inputs, and $\star$ denotes 2D pose heatmap inputs. Best results are in bold, while second best ones are underlined.
0.A.7 Qualitative Results
@@
0.A.8 Discussions
Limitations. There exist potential limitations of our 2D skeleton heatmap-based approach, including missing depth and context details. In contrast to 3D skeleton inputs (e.g., CASA [41]), 2D skeleton inputs do not contain depth information. Therefore, our method may suffer in scenarios where depth cues are crucial in classifying the actions, e.g., occlusions and extreme viewpoint differences. However, we would like to note that depth cues can still likely be inferred implicitly from 2D skeleton heatmaps, in a similar manner as monocular depth prediction, and it is challenging to predict depths accurately in cases of occlusions and extreme viewpoint differences. Furthermore, both 3D skeleton inputs (e.g., CASA [41]) and 2D skeleton inputs do not include context details, e.g., objects and background information, which are available in RGB video inputs. Thus, our method may fail in cases where context details are important in recognizing the actions. Nevertheless, our multi-modality fusion method can overcome this limitation by utilizing both 2D skeleton heatmaps and RGB videos as inputs.
Societal Impacts. Our sequence alignment method offers valuable applications including frontline worker training and assistance. In particular, frontline workers could benefit from guidance provided by models generated from expert demonstration videos in various domains from factory assembly to medical surgery. However, we also acknowledge the potential for misuse in surveillance and monitoring of individuals. This may raise privacy concerns, emphasizing the importance of responsible AI principles to guide the use of this technology.

### v6

@@ 4.1.3 Robustness against Missing Keypoints
To assess the robustness of our method, we perform an experiment where a certain part of the skeleton is removed. Specifically, we randomly remove one finger with a probability $p$ for each frame in the H2O dataset during testing (without any finetuning) and measure the impact of this perturbation on the classification accuracy. The results in Tab. 3 show that LA2DS exhibits a high level of robustness to missing keypoints. For example, when one finger is dropped per frame (i.e., $p=100\%$), the decrease in accuracy for LA2DS is minimal (about $1\%$), whereas for CASA [43], the drop is around $8\%$. This is likely because our method represents each joint by a Gaussian distribution and performs self-attention in the spatial domain to leverage dependencies between joints in the same skeleton, whereas CASA models each joint merely by its coordinates and does not conduct spatial self-attention.
@@ 4.1.4 Robustness against Noisy Keypoints
To examine the generalization of our approach to noisy keypoints, we train and test on skeletons of different qualities. In particular, we use ground truth keypoints provided by the H2O dataset, namely clean keypoints, as well as add random noise to them, yielding noisy keypoints. We measure the impact of this perturbation on the classification accuracy. The results in Tab. 4 show that LA2DS exhibits a high level of robustness against noisy keypoints. For instance, when noisy keypoints are used for both training and testing, the accuracy of CASA [43] is decreased by more than 5%, whereas our method has a drop of less than 2%. Similar to the prior section, our method is more robust to noisy keypoints than CASA, thanks to our heatmap representation and self-attention in the spatial domain.
@@ 4.2 Comparison Results
Below we quantitatively compare our approach with previous methods. In addition, some qualitative comparisons are provided in the supplementary material.
@@ 4.3 Fusion Results
4.4 Comparisons with Recent RGB-Based Methods
We compare our approach with recent RGB-based methods, namely VSP [80] (video inputs) and STEPs [63] (video and optical flow inputs). These methods do not employ temporal alignment losses. The results on IKEA ASM are shown in Tab. 9. Our method outperforms VSP and STEPs on very few metrics, i.e., better retrieval results but worse classification results than VSP, and better Acc@0.1 but worse Acc@0.5 and Acc@1.0 than STEPs. Our worse performance is likely due to our compact inputs (e.g., missing context details). Nevertheless, the key advantage of our method is privacy preserving, which is important for several applications involving human action understanding.
Method Input Acc@0.1 Acc@0.5 Acc@1.0 AP@5 AP@10 AP@15
IKEA ASM VSP [80] $\square$ 45.00 45.97 47.52 30.23 - -
STEPs [63] $\approx$ + $\square$ 28.59 36.25 37.02 - - -
LA2DS (Ours) $\star$ 26.43 32.56 34.73 32.44 31.89 31.56
LA2DS (Ours) $\star$ + $\square$ 31.74 34.51 36.18 33.08 32.25 32.17
Table 9: Comparisons with recent RGB-based methods. Note that $\square$ denotes direct RGB inputs, $\approx$ denotes optical flow inputs, and $\star$ denotes 2D pose heatmap inputs. Best results are in bold, while second best ones are underlined.
Supplementary Material. We provide additional results and discussions in the supplementary material. They include our results with different method parts, which analyze sources of performance gains, our results with OpenPose [8] 2D poses, which are better than our above results with projected 2D poses, and discussions of limitations.
@@ 5 Conclusion
We present a 2D skeleton-based self-supervised temporal video alignment framework. Specifically, we propose to take 2D skeleton heatmap inputs and employ video transformers to perform self-attention both in spatial and temporal domains. This is in contrast with CASA, which processes 3D skeleton coordinates and does not conduct spatial self-attention. For self-supervised learning, we exploit simple heatmap augmentation techniques based on 2D skeletons. Our approach achieves both better accuracy and higher robustness than CASA and shows superior results on fine-grained human activity understanding tasks on Penn Action, IKEA ASM, and H2O. Finally, our multi-modality version establishes the state-of-the-art on all metrics and datasets. To our best knowledge, our work is the first to exploit 2D skeleton heatmap inputs and the first to explore multi-modality fusion for temporal video alignment.
@@ Appendix 0.A Supplementary Material
In this supplementary material, we first provide the details of our implementation in Sec. 0.A.1. Next, the ablation results of different augmentations are included in Sec. 0.A.2, while our multi-modality model architecture and our complete multi-modality fusion results on all metrics and datasets are presented in Sec. 0.A.3. We then report the ablation results of different method parts in Sec. 0.A.4, while the ablation results of different 2D poses are included in Sec. 0.A.5. Lastly, we show some qualitative results in Sec. 0.A.6 and discuss the limitations and societal impacts of our work in Sec. 0.A.7.
@@ 0.A.3 Multi-Modality Fusion
We first illustrate our multi-modality model architecture in Fig. 6, which we describe in details in Sec. 3.4 of the main paper. Next, our complete multi-modality fusion results on all metrics and datasets are presented in Tab. 12. It is clear from the results, our RGB+Pose model yields the best numbers across all metrics and datasets, outperforming all the single-modality methods and the multi-modality baseline extended from CASA [43] and establishing the state-of-the-art across all metrics and datasets. In addition, our Pose-only model obtains better results than our RGB-only model on Penn Action and H2O, while having generally worse results on IKEA ASM. It likely implies that pose cues are more important on Penn Action and H2O, while RGB cues are more crucial on IKEA ASM. Finally, both our RGB+Pose model and the CASA-based multi-modality baseline outperform their single-modality counterparts respectively, which validates the benefit of multi-modality inputs for fine-grained human activity understanding tasks.
@@ 0.A.5 Effects of Different 2D Poses
0.A.6 Qualitative Results
@@
0.A.7 Discussions
0.A.8 Limitations
Challenging Viewpoints and Extreme Viewpoint Gaps. We follow all of previous works to use Penn Action, IKEA ASM, and H2O, which contain frontal, side, or top viewpoints of human actions (e.g., Fig. 5 of the main paper), since it is difficult to capture human actions with rear or bottom viewpoints due to severe occlusions or privacy concerns respectively. Also, it is generally harder to obtain accurate 3D poses than accurate 2D poses, especially for challenging viewpoints, thus CASA [43]’s sensitivity to noise/errors of pose inputs may further be amplified with challenging viewpoints. However, we do randomly pair videos to perform alignment, yielding viewpoint gaps of up to $180$ degrees (e.g., pairing videos with left and right side viewpoints in Figs. 5(a-b) of the main paper). Collecting a human action dataset including rear or bottom viewpoints and learning by aligning videos with extreme viewpoint gaps are challenging and remain our future works.
Missing Context Details. Both 3D skeleton inputs (e.g., CASA [43]) and 2D skeleton inputs (i.e., our method) do not include context details, e.g., objects and background information, which are available in RGB video inputs. Thus, our method may fail in cases where context details are important in recognizing the actions. Nevertheless, our multi-modality fusion method can overcome this limitation by utilizing both 2D skeleton heatmaps and RGB videos as inputs.
0.A.9 Societal Impacts
Our sequence alignment method offers valuable applications including frontline worker training and assistance. In particular, frontline workers could benefit from guidance provided by models generated from expert demonstration videos in various domains from factory assembly to medical surgery. However, we also acknowledge the potential for misuse in surveillance and monitoring of individuals. This may raise privacy concerns, emphasizing the importance of responsible AI principles to guide the use of this technology.